OpenStudy (anonymous):

two angles are complementary. the sum of the measure of the first angle and one-fourth the second angle is 60.75 degrees find the measures of the angles. what is the measure of the smaller angle? what is the measure of the other angle?

OpenStudy (anonymous):

x+y=90 x+y/4=60.75

OpenStudy (anonymous):

2 angles are complementary if they add up to 90 degree. Thus you can set up the equations to: angle 1 + angle2 =90 angle 1 + 1/4 angle 2 = 60.75 You can find method to solve system of linear equation at: http://www.oojih.com/show/algebra/syslinearequation/ check out www.oojih.com for free study guide and free review questions.

OpenStudy (anonymous):

(1-1/4)y=29.25 3y/4=29.25 y=29.25*4/3=39 x=51

OpenStudy (anonymous):

uzma so smaller angle is 39 and other angle is 51 correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thansk uzma for all the info

OpenStudy (anonymous):

your welcome dear